OpenStudy (anonymous):

Solve for c. ab + c = 5 c = 5/ab c = 5 - ab c = 5 + ab

Parth (parthkohli):

c = 5 - ab

Parth (parthkohli):

I got it by transposing.

jimthompson5910 (jim_thompson5910):

you can rewrite the left side of ab+c to get c+ab So we're adding 'ab' to 'c', to undo this, we subtract 'ab' from both sides c+ab - ab = 5 - ab c + 0ab = 5 - ab c + 0 = 5 - ab c = 5 - ab We did this so the 'ab' terms on the left side will add to 0 and cancel out.